A Bibliography of Publications of Anders Forsgren

Anders Forsgren

Royal Institute of Technology (KTH)

S-100 44 Stockholm Sweden

Tel: +46 8 790 71 27 FAX: +46 8 22 53 20

E-mail: andersf@math.kth.se (Internet)

13 October 2017 Version 1.10

Abstract

This bibliography records publications of Anders Forsgren.

Title word cross-reference

16 [For08].

algorithm [BF13]. algorithms [For02]. analysis [FR92, FR93, OHCF16]. approaches [CF14]. approximate [CF14]. approximating [BF13]. Arising [FGG07, FGS94, FGS96]. Augmented [FGG07].

based [BF13]. between [FO15].

Cholesky [FGM93, FGM95]. column [CF14, OHCF16]. Comments [For08]. communications [PF02]. Computing [FGM93, FGM95]. conditioned [FGS94, FGS96]. conditions [For00]. conjugate [FO15]. connection [FO15]. Constrained [FM93, FM97, FM90, FGS94, FGS96]. Controlling [FGM91, FGM89b, For02]. Convex [FS01, BF13].

Dempsey [For08]. Diagonally [For96]. Dimensioning [PF02]. Directions [FGM95, FGM93]. Dominant [For96]. Dual [FG98, BF13]. dynamically [OHCF15].

element [FR92, FR93]. elementary [OHCF15, OHCF16]. enabled [PF02]. Equality [FM93, FM90]. Equality-Constrained [FM93, FM90].

F [For08]. Factorization [FGM95, FGM93]. factorizations [For02]. finite [FR92, FR93]. flux [OHCF15, OHCF16].

generated [OHCF16]. generating [OHCF15]. generation [CF14, OHCF16]. gradient [FO15].

H [For08].

Identification [FGM91, FGM89b]. ill [FGS94, FGS96]. ill-conditioned [FGS94, FGS96]. Inequality [FM97]. Inequality-Constrained [FM97]. Inertia [FGM91, For02, FGM89b]. Inertia-Controlling [FGM91, For02, FGM89b]. Intensity [For08, CF14]. intensity-modulated [CF14]. Interior [FG98, FS01, FGW02, FGG07, FGS94, FGS96, For06]. Iterative [FGG07].

J [For08].

Large [FM93, FM97, FM90]. Large-Scale [FM93, FM97, FM90]. Least [For96, FS01]. Least-Squares [For96, FS01]. Linear [FM93, For96, FM97, FS01, FM90]. Local [FGM91, FGM89b].

Matrices [For96]. metabolic [OHCF15]. method [FGM89a, FR92, FR93, FO15]. Methods [FGM91, FM93, FM97, FG98, FS01, FGW02, FGG07, FGM89b, FM90, For90, FGS94, FGS96, For06, FO15]. Minimization [FM93, FM97, FGM89a, FM90]. Minimizers [FGM91, FGM89b]. modes [OHCF15, OHCF16]. Modified [FGM95, FGM89a, FR92, FGM93, FR93]. modulated [CF14, For08]. multicast [PF02]. multicast-enabled [PF02].

network [OHCF15]. networks [PF02]. Newton [FGM89a, FM90, For90, FR92, FGM93, FM93, FGM95, FM97, FO15, FR93]. no [For08]. Nonconvex [FG98, For90, For00]. Nonlinear [FG98, FGW02, FR92, FR93].

Optimality [For00]. Optimization [FGW02, For90, FGS94, FGS96, For02, For08, OHCF15].

Pareto [BF13]. Partial [FGM95, FGM93]. plan [For08]. Primal [FG98]. Primal-Dual [FG98]. Problems [For96, FS01, FO15]. Programming [FGM91, FG98, FS01, FGM89b, For00]. programs [CF14].

Quadratic

[FGM91, FS01, CF14, FGM89b, FO15]. quasi [FO15]. quasi-Newton [FO15].

radiation [CF14, For08]. Related [FS01]. relevant [OHCF15]. Robustness [OHCF16]. Romeijn [For08].

Scale [FM93, FM97, FM90]. semidefinite [For00]. Solution [FGG07]. solutions [CF14]. Squares [For96, FS01]. Stability [FGS94, FGS96]. starts [For06]. surfaces [BF13]. symmetric [FGS94, FGS96]. Systems [FGG07, FGS94, FGS96].

techniques [BF13]. therapy [CF14, For08]. TOP [For08]. treatment [For08].

unconstrained [FGM89a]. use

[FR92, FR93]. Using

[FGM95, FGM93, OHCF15].

warm [For06]. Weight [For96]. Weighted [FS01].

References

Bokrantz:2013:AAC

[BF13] Rasmus Bokrantz and Anders Forsgren. An algorithm for approximating convex Pareto surfaces based on dual techniques. IN-FORMS Journal on Computing, 25 (2):377–393, 2013. ISSN 1091-9856 (print), 1526-5528 (electronic).

Carlsson:2014:CGA

[CF14] Fredrik Carlsson and Anders Forsgren. On column generation approaches for approximate solutions of quadratic programs in intensity-modulated radiation therapy. Annals of Operations Research, 223:471–481, 2014. CO-DEN AOREEV. ISSN 0254-5330 (print), 1572-9338 (electronic).

Forsgren:1998:PDI

[FG98] Anders Forsgren and Philip E.

Gill. Primal-dual interior methods for nonconvex nonlinear programming. SIAM Journal on Optimization, 8(4):1132–1152, November 1998. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic). URL http:/

/epubs.siam.org/sam-bin/dbq/
article/30556.

Forsgren:2007:ISA

[FGG07] Anders Forsgren, Philip E. Gill, and Joshua D. Griffin. Iterative solution of augmented systems arising in interior methods. SIAM Journal on Optimization, 18(2): 666–690, ???? 2007. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Forsgren:1989:MNM

[FGM89a] A. Forsgren, P. E. Gill, and W. Murray. A modified Newton method for unconstrained minimization. Report SOL 89-12, Department of Operations Research, Stanford University, 1989.

Forsgren:1989:ILM

[FGM89b] A. Forsgren, P. E. Gill, and W. Murray. On the identification of local minimizers in inertiacontrolling methods for quadratic programming. Report SOL 89-11, Department of Operations Research, Stanford University, 1989.

Forsgren:1991:ILM

[FGM91] A. L. Forsgren, P. E. Gill, and W. Murray. On the identification of local minimizers in inertiacontrolling methods for quadratic programming. SIAM Journal on Matrix Analysis and Applications, 12(4):730–746, October 1991. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Forsgren:1993:CMN

[FGM93] A. Forsgren, P. E. Gill, and W. Murray. Computing modified Newton directions using a partial Cholesky factorization. Report TRITA-MAT-1993-9, Division of Optimization and Systems Theory, Department of Mathematics, Royal Institute of Technology, 1993.

Forsgren:1995:CMN

[FGM95] Anders Forsgren, Philip E. Gill, and Walter Murray. Computing modified Newton directions using a partial Cholesky factorization. SIAM Journal on Scientific Computing, 16(1):139–150, January 1995. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Forsgren:1994:SSI

[FGS94] A. Forsgren, P. E. Gill, and J. R.

Shinnerl. Stability of symmetric illconditioned systems arising in interior methods for constrained optimization. Report TRITA-MAT-1994-24, Division of Optimization and Systems Theory, Department of Mathematics, Royal Institute of Technology, 1994. Published in [FGS96].

Forsgren:1996:SSI

[FGS96] Anders Forsgren, Philip E. Gill, and Joseph R. Shinnerl. Stability of symmetric ill-conditioned systems arising in interior methods for constrained optimization. SIAM Journal on Matrix Analysis and Applications, 17(1):187–211, January 1996. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). URL http:/

/epubs.siam.org/sam-bin/dbq/
article/27065.

Forsgren:2002:IMN

[FGW02] Anders Forsgren, Philip E. Gill, and Margaret H. Wright. Interior methods for nonlinear optimization. SIAM Review, 44(4): 525–597, December 2002. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/41494.

Forsgren:1990:NML

[FM90] A. Forsgren and W. Murray. Newton methods for large-scale linear equality-constrained minimization. Report SOL 90-6, Department of Operations Research, Stanford University, 1990.

Forsgren:1993:NML

[FM93] A. Forsgren and W. Murray. Newton methods for large-scale linear equality-constrained minimization. SIAM Journal on Matrix Analysis and Applications, 14(2): 560–587, April 1993. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Forsgren:1997:NML

[FM97] Anders Forsgren and Walter Murray. Newton methods for largescale linear inequality-constrained minimization. SIAM Journal on Optimization, 7(1):162–176, February 1997. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/dbq/
                                        article/27912.

Forsgren:2015:CBC

[FO15] Anders Forsgren and Tove Odland. On the connection between the conjugate gradient method and quasi-Newton methods on quadratic problems. Computational Optimization and Applications, 60(2):377–392, 2015. CO-DEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic).

Forsgren:1990:NMN

[For90] A. Forsgren. Newton methods for nonconvex optimization. PhD thesis, Report TRITA-MAT-1990-10, Division of Optimization and Systems Theory, Department of Mathematics, Royal Institute of Technology, Stockholm, Sweden, 1990.

Forsgren:1996:LLS

[For96] Anders Forsgren. On linear leastsquares problems with diagonally dominant weight matrices. SIAM Journal on Matrix Analysis and Applications, 17(4):763–788, October 1996. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). URL http:/

/epubs.siam.org/sam-bin/dbq/
article/28401.

Forsgren:2000:OCN

[For00] Anders Forsgren. Optimality conditions for nonconvex semidefinite programming. Mathematical Programming, 88(1, Ser. A):105–128, 2000. CODEN MHPGA4. ISSN 0025-5610.

Forsgren:2002:ICF

[For02] Anders Forsgren. Inertia-controlling factorizations for optimization algorithms. Applied Numerical Mathematics: Transactions of IMACS, 43(1–2):91–107, October 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). 19th Dundee Biennial Conference on Numerical Analysis (2001).

Forsgren:2006:WSI

[For06] A. Forsgren. On warm starts for interior methods. In System modeling and optimization, volume 199 of IFIP Int. Fed. Inf. Process., pages 51–66. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006.

Forsgren:2008:CIM

[For08] A. Forsgren. Comments on: “Intensity modulated radiation therapy treatment plan optimization” [TOP 16 (2008), no. 2, 215–243; Cno. 2453404] by H. E. Romeijn and J. F. Dempsey. TOP, 16(2): 246–247, 2008. ISSN 1134-5764.

Forsgren:1992:UMN

[FR92] A. Forsgren and U. Ringertz. On the use of a modified Newton method for nonlinear finite element analysis. Report TRITA-MAT-1992-22, Division of Optimization and Systems Theory, Department of Mathematics, Royal Institute of Technology, 1992.

Forsgren:1993:UMN

[FR93] Anders Forsgren and Ulf Ringertz. On the use of a modified Newton method for nonlinear finite element analysis. Computer Methods in Applied Mechanics and Engineering, 110(3-4):275–283, 1993. CODEN CMMECC. ISSN 0045-7825, 0374-2830.

Forsgren:2001:WLL

[FS01] Anders Forsgren and Göran Sporre. On weighted linear leastsquares problems related to interior methods for convex quadratic programming. SIAM Journal on Matrix Analysis and Applications, 23(1):42–56, 2001. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/37229.

Oddsdottir:2015:DGR

[OHCF15] Hildur Æsa Oddsdóttir, Erika Hagrot, Véronique Chotteau, and Anders Forsgren. On dynamically generating relevant elementary flux modes in a metabolic network using optimization. Journal of Mathematical Biology, 71(4): 903–920, October 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:

//link.springer.com/content/
pdf/10.1007/s00285-014-0844-
1.pdf.

Oddsdottir:2016:RAE

[OHCF16] Hildur Æsa Oddsdóttir, Erika Hagrot, Véronique Chotteau, and Anders Forsgren. Robustness analysis of elementary flux modes generated by column generation. Mathematical Biosciences, 273:45–56, 2016. CODEN MABIAR. ISSN 0025-5564 (print), 1879-3134 (electronic).

Prytz:2002:DME

[PF02] Mikael Prytz and Anders Forsgren.

Dimensioning multicast-enabled communications networks. Networks, 39(4):216–231, 2002. CO-DEN NTWKAA. ISSN 0028-3045 (print), 1097-0037 (electronic).